how much of a 20% solution of nitric acid should be added to 3 L of 12% solution to produce a 15% solution?
Answer by mananth(16946)

--------percent ---------------- quantity
Nitric acid I 12 ---------------- 3 L
Nitric Acid II 20 ---------------- x L
Mixture 15 ---------------- 3 + x L

12 * 3 + 20 x = 15 ( 3 + x )
36 + 20 x = 45 + 15 x
20 x + -15 x = 45 - -36
5 x = 9
/ 5
x = 1.8
1.8 L of Nitric Acid II